# Probabilistic Time Integration for Semi-explicit PDAEs

### v1

@@ 1. Introduction
Over the past few years, randomized time integration methods for ordinary differential equations have been studied extensively; see [CCCG16, CGS+17, LSS19, AG20, LSS22]. These methods aim to statistically quantify the uncertainty introduced by the time discretization. For this, classical probabilistic solvers iteratively establish a probability measure over the numerical solution of deterministic initial value problems of the form
(1.1) $\dot{u}(t)=f(t,u(t)),\qquad u(0)=u^{0}\in\mathbb{R}^{d},$
offering more comprehensive information than a single (deterministic) solve. Let $\psi_{t}\colon[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denote the flow map induced by (1.1), meaning that $u(t)=\psi_{t}(0,u^{0})$. On the other hand, a numerical one-step method can be expressed by a mapping $\Psi_{t}\colon[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with
$u^{n+1}=\Psi_{\tau}(t^{n},u^{n}),\qquad n=0,1,\dots,$
assuming an equidistant partition with $t^{n}=n\tau$ and time step size $\tau$. Here, $u^{n}$ denotes the approximation of $u(t^{n})=\psi_{n\tau}(t^{0},u^{0})$. Typically, these methods produce a single discrete solution, often accompanied by some form of an error indicator. However, they do not fully quantify the remaining uncertainty in the path statistically. In order to capture the sensitivity of the solution to numerical errors, while maintaining consistent convergence properties from classical deterministic integrators, the probabilistic interpretation of the numerical solution introduced in [CGS+17] considers the randomized scheme
@@
with $U^{0}=u^{0}$ and appropriately scaled independent and identically distributed (i.i.d.) random variables $\xi^{n}(\tau)$ with values in $\mathbb{R}^{d}$. This then results in a sequence of random variables $U^{n}$ approximating $u(t^{n})$. We would like to emphasize that probabilistic integrators may neither inherently provide more accurate solutions than classical deterministic methods, nor are they necessarily computationally cheaper. They prove useful, however, in statistical inference settings, particularly in tasks such as integrating chaotic dynamical systems and solving Bayesian inverse problems [CCCG16].
@@
Known deterministic time stepping methods for such DAEs and PDAEs include algebraically stable Runge–Kutta methods [HLR89, KM06, AZ18], splitting methods [AO17], (discontinuous) Galerkin methods [VR19, AH21], and exponential integrators [AZ20]. In this study, we provide a framework how to randomize existing (deterministic) integration schemes, including the specific construction of four randomized methods. We identify assumptions and introduce a local random field, particularly a Gaussian field, to reflect the uncertainty. The resulting probabilistic solvers allow for repeated sampling in order to explore the solutions uncertainty. In the following example, we illustrate the difficulty arising for constrained systems and the consequences of perturbations affecting the constraint; see also [ALM17].
@@
To illustrate the randomized idea from [CGS+17], let us consider the FitzHugh–Nagumo model [RHCC07] with an additional constraint. With the help of the Lagrangian method, we obtain the nonlinear system
@@
with the constraint $V(t)+R(t)=\sin(t)$ and unknowns $V$, $R$, and $\lambda$. As initial conditions, we set $V(0)=-1$ and $R(0)=1$. When a small perturbation is added to the constraint, implemented as a scalar Gaussian random variable $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, this leads to chaotic behavior. As a result, the trajectories strongly deviate from the true solution. However, if we add perturbations only in the dynamic part, one can see that every probabilistic solution is a good approximation of the exact solution. In Figure 1.1, we present $50$ trajectories for the $V$ species computed by the implicit Euler scheme with $\sigma=0.1$, once affecting the constraint and once only acting on the dynamic part of the solution. This illustrates the sensitivity of the solution to perturbations on the constraint. On the other hand, Figure 1.2 depicts $50$ trajectories for the $V$ species for different noise scales $\sigma$, showing the sensitivity of the solutions with respect to $\sigma$. At this point, we would like to emphasize that the scale parameter $\sigma$ should, in general, be chosen problem-dependent on basis of specific characteristics of the model. Such a calibration controls the apparent uncertainty in the solver and is discussed later on.
@@
In the forthcoming section, we introduce the model of interest and the necessary setup for the construction of probabilistic integrators. The mean square convergence is then analyzed in Section 3. In Section 4, we delve into the construction of four specific integrators, namely the implicit Euler method, the midpoint scheme, and two exponential integrators of first and second order. To confirm the theoretical findings, we provide a numerical example. Finally, Section 5 is devoted to the calibration of the parameter $\sigma$ when employing randomized time integrators for constrained parabolic systems. Here, we demonstrate that an appropriately scaling enables the randomized solvers to yield probabilistic solutions without affecting the convergence of the corresponding deterministic scheme.
Notation
Given $N\in\mathbb{N}$, we define $[N]\coloneqq\{1,\ldots,N\}$ as well as $[N]_{0}\coloneqq[N]\cup\{0\}$. Moreover, given discrete time points $t^{n}$ and a (continuous) function $g$, we write $g^{n}\coloneqq g(t^{n})$.
@@ 2.1. Semi-explicit (P)DAEs
Within this section, we introduce the model of interest, namely (semi-linear) parabolic systems with linear constraints. Note that this is a PDAE, since the corresponding semi-discrete system, which is obtained by the application of a spatial discretization, equals a semi-explicit DAE of index $2$. More precisely, we consider following task: seek abstract functions $u\colon[0,T]\to\mathcal{V}$ and $\lambda\colon[0,T]\to\mathcal{Q}$ such that
@@
The differential operator $\mathcal{A}\colon\mathcal{V}\to\mathcal{V}^{*}$ is linear, continuous, and elliptic on $\mathcal{V}_{\ker}\coloneqq\ker\mathcal{B}$, i.e., on the kernel of the constraint operator. Restricted to the kernel, we denote the operator by $\mathcal{A}_{\ker}\coloneqq\mathcal{A}|_{\mathcal{V}_{\ker}}\colon\mathcal{V}_{\ker}\to\mathcal{V}_{\ker}^{*}\coloneqq(\mathcal{V}_{\ker})^{*}$.
@@
The upcoming results can be extended to the situation where $\mathcal{A}$ only satisfies a Gårding inequality on $\mathcal{V}_{\ker}$. In this case, the non-elliptic part can be transferred to the nonlinearity $f$.
@@
The right-hand sides are assmued to be sufficiently smooth to guarantee the existence of a unique solution. In particular, we assume that $f$ maps into $\mathcal{H}$ and that it is Lipschitz continuous in the second component, i.e., there exists a positive constant $L_{f}$ such that
@@ 2.2. Decomposition of the solution and flow map
The already mentioned consistency condition of the initial data can also be understood in the way that we are only allowed to prescribe $u_{\ker}(0)$. The remaining part $u_{c}(0)$ is determined by $\mathcal{B}^{-}g(0)$ and $\lambda(0)$ can be computed by (2.1a). This motivates the definition of a flow map on the kernel, which we denote by $\psi_{t}\colon[0,T]\times\mathcal{V}_{\ker}\rightarrow\mathcal{V}_{\ker}$. For given $t^{*}\in[0,T]$ with $t^{*}+t\leq T$ and $v_{\ker}\in\mathcal{V}_{\ker}$, this map solves the PDAE (2.1) on the time interval $[t^{*},t^{*}+t]$ with initial data $v_{\ker}+\mathcal{B}^{-}g(t^{*})$. The outcome is then the solution at time $t^{*}+t$ restricted to kernel of $\mathcal{B}$. In particular, we have
@@ 2.3. Deterministic and probabilistic setup
By $\Psi_{\tau}$ we denote a (deterministic) time integration method which can be used to approximate the solution of the semi-explicit system (2.1) or its semi-discretized counterpart. Based on such a scheme, we define a randomized integrator, resulting in a sequence $\{U^{n}\}_{n\in[N]_{0}}$. Each approximation can again be decomposed into
$U^{n}=U^{n}_{\ker}+U^{n}_{c},\qquad U^{n}_{\ker}\in\mathcal{V}_{\ker},\ U^{n}_{c}\in\mathcal{V}_{\text{c}}.$
As for the continuous solution, we assume that the approximation at time $t^{n+1}$ is fully determined by $U^{n}_{\ker}$ and the right-hand sides. Hence, we write
@@
where each $\xi^{n}$ is assumed to be an i.i.d. centred Gaussian random variable with values in $\mathcal{V}_{\ker}$. Since the perturbation only acts on the dynamic part of the solution, the constraint $\mathcal{B}U^{n}=g(t^{n})$ is maintained for all discrete time points. The sequence of errors $e^{n}$ of the random approximation is defined by
@@
The assumptions on the (deterministic) scheme $\Psi_{\tau}$ are summarized in the following.
@@
The i.i.d. centred Gaussian random variables $\xi^{n}(t)\coloneqq\int_{0}^{t}\chi^{n}(s)\,\mathrm{d}s\in\mathcal{V}_{\ker}$ with $\chi^{n}\sim\mathcal{N}(0,\sigma^{2})$ measured in the $\mathcal{H}$-norm, i.e.,
$\mathbb{E}\,\|\chi^{n}\|_{\mathcal{H}}=0,\qquad\mathbb{E}\,\|\chi^{n}\|^{2}_{\mathcal{H}}=\sigma^{2},$
admit parameters $C_{\xi},p>0$ such that for all $t\in[0,\tau]$ it holds that $\mathbb{E}\,\|\xi^{n}(t)\|_{\mathcal{H}}^{2}\leq C_{\xi}\,t^{2p+1}$.
Within the upcoming convergence analysis, we will observe that choosing $p=q$ preserves the strong order of accuracy of the underlying deterministic integrator. This choice introduces the maximum amount of noise consistent with the accuracy of the original deterministic integrator.
@@ 3. Mean Square Convergence Analysis
for all $n\in[N]$. Therein, $r$ is the mean square order of convergence of the method. For the unconstrained case, it is known that $r$ equals to the minimum of $q$ (convergence order of the deterministic time integrator) and $p$ (order of the random noise). Given that our approach is based on a construction in the kernel of the constraint, we expect the same order for the given setting. The analysis of this section will make use of the following inequality.
Lemma 3.1 (Discrete Gronwall lemma).
@@
We can now state the global mean square convergence result for probabilistic time integrators applied to (2.1).
@@
where Assumption 2.7 was applied in the last step. Now we consider the expectation of this estimate. Since $\xi^{n}$ equals a zero-mean Gaussian random variable, the expectation of the inner product vanishes and we get together with Assumption 2.8
@@
Finally, an application of the triangle inequality
$\mathbb{E}\,\big{\|}e^{n}\big{\|}^{2}_{\mathcal{H}}$ $\leq 2\,\mathbb{E}\,\big{\|}\tilde{e}^{n}\big{\|}^{2}_{\mathcal{H}}+2\,\mathbb{E}\,\big{\|}e^{n}-\tilde{e}^{n}\big{\|}^{2}_{\mathcal{H}},$
Young’s inequality, and the assumption on the global error of the deterministic solver given in (2.7) yields the desired result. ∎
@@
Condition (2.7) on the order of the deterministic solver may also be replaced by an assumption on the local truncation error. If one assumes a local error of order $q+1$, mean square order of convergence will still be $\min\{p,q\}$. However, the difference between the global and local order of convergence, does not follow the typical pattern observed in deterministic cases. Indeed, the independence of the random variables results in an order reduction of only $1/2$ in the random part of the exponent, while the expected loss of a full order occurs in the deterministic component.
Remark 3.4.
The convergence result of Theorem 3.2 implies that a reasonable option for the noise scale $p$ is to set $p=q$, where $q$ is the order of the underlying deterministic method. This choice of $p$ preserves the convergence of the underlying deterministic method, while providing a probabilistic interpretation of the numerical solution.
@@ 4.1. Probabilistic implicit Euler scheme
(4.1b) $\mathcal{B}u^{n+1}\hskip 80.52135pt$ $=\phantom{\tau}g^{n+1}\hskip 41.82558pt\qquad\text{in }\mathcal{Q}^{*}.$
Before incorporating perturbations, we discuss Assumption 2.7. It is well-known that the implicit Euler scheme converges with order one without constraints [LO93]. This then directly translates to the solution part in the kernel of the constraint, i.e., we get (2.7) with $q=1$. To prove the Lipschitz property (2.8), we use once more the decomposition of the iterates $u^{n}=u^{n}_{\ker}+u^{n}_{c}$ with $u^{n}_{c}=\mathcal{B}^{-}g^{n}$ as described in Section 2.2. Restricting equation (4.1a) to test functions in $\mathcal{V}_{\ker}$, we get
$u^{n+1}_{\ker}=u^{n}_{\ker}+\tau\,\big{(}-\mathcal{A}\,u^{n+1}_{\ker}+f(t^{n+1},u_{\ker}^{n+1}+\mathcal{B}^{-}g^{n+1})\big{)}-\mathcal{B}^{-}\big{(}g^{n+1}-g^{n}\big{)}\quad\text{in }\mathcal{V}_{\ker}^{*}$
and, hence,
@@
(4.2) $\begin{bmatrix}\operatorname{id}+\tau\mathcal{A}&\tau\mathcal{B}^{*}\\ \mathcal{B}&0\end{bmatrix}\begin{bmatrix}U^{n+1}\\ \lambda^{n+1}\end{bmatrix}=\begin{bmatrix}U^{n}+\tau f(t^{n+1},U^{n+1})+\xi^{n}(\tau)\\ g^{n+1}\end{bmatrix},$
where $\xi^{n}(\tau)\in\mathcal{V}_{\ker}$ can be defined as $\sigma\tau^{3/2}\tilde{\xi}^{n}$ by the natural choice of $p=q=1$. Here, $\sigma$ is a constant noise scale and $\tilde{\xi}^{n}$ are i.i.d. standard normal random variables i.e., $\tilde{\xi}^{n}\sim\mathcal{N}(0,1)$. In Section 5, we discuss the case where $\sigma$ is chosen problem-dependent such that the expected order of convergence is maintained.

### v3

@@ Abstract.
This article will be published in Statistics and Computing.
@@ 1. Introduction
In the last decade, randomized time integration methods for ordinary, partial, and operator differential equations have been studied extensively; see [CCCG16, CGS+17, LSS19, AG20, LSS22]. These methods aim to statistically quantify the uncertainty introduced by the time discretization. For this, classical probabilistic solvers iteratively establish a probability measure over the numerical solution of deterministic initial value problems of the form
$\dot{u}(t)=$ $f(t,u(t)),\qquad\text{for $0\leq t\leq T$,}$
$u(0)=$ $u^{0}\in\mathbb{R}^{d},$
offering more comprehensive information than a single (deterministic) solve. This means that, instead of yielding just one solution path, the probabilistic approach offers a distribution over possible solutions. Consequently, such methods capture the uncertainty inherent in the numerical solution process, especially in regions where the solution is not directly computed. Let $\psi_{t}\colon[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denote the flow map induced by (1), meaning that $u(t)=\psi_{t}(0,u^{0})$. On the other hand, we consider a discretization scheme based on a (constant) time step size $\tau>0$ with $N\coloneqq T/\tau\in\mathbb{N}$ and corresponding time points $t^{n}\coloneqq n\tau$ for $n\in[N]\coloneqq\{0,1,\ldots,N\}$. With this, a numerical one-step method can be expressed by a mapping $\Psi_{t}\colon[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with
$u^{n+1}=\Psi_{\tau}(t^{n},u^{n}),\qquad n\in[N].$
Here, $u^{n}$ denotes the approximation of $u(t^{n})=\psi_{n\tau}(t^{0},u^{0})$. Typically, these methods produce a single discrete solution, often accompanied by some form of an error indicator. However, they do not fully quantify the remaining uncertainty in the path statistically. In order to capture the sensitivity of the solution to numerical errors, while maintaining consistent convergence properties from classical deterministic integrators, the probabilistic interpretation of the numerical solution introduced in [CGS+17] considers the randomized scheme
@@
with $U^{0}=u^{0}$ and appropriately scaled independent and identically distributed (i.i.d.) random variables $\xi^{n}(\tau)$ with values in $\mathbb{R}^{d}$. This then results in a sequence of random variables $U^{n}$ approximating $u(t^{n})$. We would like to emphasize that probabilistic integrators may neither inherently provide more accurate solutions than classical deterministic methods, nor are they necessarily computationally cheaper. They prove useful, however, in statistical inference settings, particularly in tasks such as integrating chaotic dynamical systems and solving Bayesian inverse problems [CCCG16, CGS+17].
@@
Known deterministic time stepping methods for such DAEs and PDAEs include algebraically stable Runge–Kutta methods [HLR89, KM06, AZ18], splitting methods [AO17], (discontinuous) Galerkin methods [VR19, AH21], and exponential integrators [AZ20]. In this study, we provide a framework how to randomize existing integration schemes, including the specific construction of four randomized methods. We identify assumptions and introduce a local random field, particularly a Gaussian field, to reflect the uncertainty of the solution in between mesh points. The resulting probabilistic solvers allow for repeated sampling in order to explore the solutions uncertainty. In the following example, we illustrate the difficulty arising for constrained systems and the consequences of perturbations affecting the constraint; see also [ALM17].
@@
To illustrate the randomized approach presented in [CGS+17], which was originally applied to the unconstrained FitzHugh–Nagumo model [RHCC07], we extend this example by an additional constraint. With the help of the Lagrangian method, we obtain the nonlinear system
@@
with the constraint $V(t)+R(t)=\sin(t)$ and unknowns $V$, $R$, and $\lambda$. As initial conditions, we set $V(0)=-1$ and $R(0)=1$. When a small perturbation is added to the constraint, implemented as a scalar Gaussian random variable $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, this leads to chaotic behavior. As a result, the trajectories strongly deviate from the true solution. However, if we add perturbations only in the dynamic part, one can see that every probabilistic solution is a good approximation of the exact solution. In Figure 1.1, we present $50$ trajectories for the $V$ species computed by the implicit Euler scheme with $\sigma=0.1$, once affecting the constraint and once only acting on the dynamic part of the solution. We refer the readers to Section 4.1 for more details on the computation of the trajectories. This illustrates the sensitivity of the solution to perturbations on the constraint. On the other hand, Figure 1.2 depicts $50$ trajectories for the $V$ species for different noise scales $\sigma$, showing the sensitivity of the solutions with respect to $\sigma$. Although we employed the same method, namely the probabilistic implicit Euler method, we varied the noise scale to demonstrate that the scale parameter $\sigma$ controls the apparent uncertainty in the solver. This variation affects the magnitude of the error induced by the probabilistic method. At this point, we would like to emphasize that the scale parameter $\sigma$ should, in general, be chosen problem-dependent on basis of specific characteristics of the model. Such a calibration controls the apparent uncertainty in the solver and is discussed later on.
@@
In the forthcoming section, we introduce the model of interest and the necessary setup for the construction of probabilistic integrators. The mean square convergence is then analyzed in Section 3. In Section 4, we delve into the construction of four specific integrators, namely the implicit Euler method, the midpoint scheme, and two exponential integrators of first and second order. To confirm the theoretical findings, we provide a numerical example. Finally, Section 5 is devoted to the calibration of the parameter $\sigma$ when employing randomized time integrators for constrained parabolic systems. Here, we demonstrate that an appropriate scaling enables the randomized solvers to yield probabilistic solutions without affecting the convergence of the corresponding deterministic scheme.
@@ 2.1. Semi-explicit (P)DAEs
Within this section, we introduce the model of interest, namely (semi-linear) parabolic systems with linear constraints. Note that this is a PDAE, since the corresponding semi-discrete system, which is obtained by the application of a spatial discretization, equals a semi-explicit DAE of index $2$. More precisely, we consider the following task: seek abstract functions $u\colon[0,T]\to\mathcal{V}$ and $\lambda\colon[0,T]\to\mathcal{Q}$ such that
@@
The differential operator $\mathcal{A}\colon\mathcal{V}\to\mathcal{V}^{*}$ is linear, continuous, and elliptic on $\mathcal{V}_{\ker}\coloneqq\ker\mathcal{B}$, i.e., on the kernel of the constraint operator. Restricted to the kernel, we denote the operator by $\mathcal{A}_{\ker}\coloneqq\mathcal{A}|_{\mathcal{V}_{\ker}}\colon\mathcal{V}_{\ker}\to\mathcal{V}_{\ker}^{*}\coloneqq(\mathcal{V}_{\ker})^{*}$. Note that we use $\mathcal{V}_{\ker}^{*}$ as codomain, which means that also the test functions are restricted to $\mathcal{V}_{\ker}$.
@@
The upcoming results can be extended to the situation where $\mathcal{A}$ only satisfies a Gårding inequality on $\mathcal{V}_{\ker}$. In this case, the non-elliptic part can be transferred to the nonlinearity $f$, cf. [AZ20].
@@
The right-hand sides are assumed to be sufficiently smooth to guarantee the existence of a unique solution. In particular, we assume that $f$ maps into $\mathcal{H}$ and that it is Lipschitz continuous in the second component, i.e., there exists a positive constant $L_{f}$ such that
@@ 2.2. Decomposition of the solution and flow map
The already mentioned consistency condition of the initial data, $\mathcal{B}u^{0}=g(0)$, determines the value of $u_{c}(0)$ but not of $u_{\ker}0)$, while $\lambda(0)$ can be computed by (2.1a). Hence, we may only prescribe a value for $u_{\ker}(0)$. This motivates the definition of a flow map on the kernel, which we denote by $\psi_{t}\colon[0,T]\times\mathcal{V}_{\ker}\rightarrow\mathcal{V}_{\ker}$. For given $t^{*}\in[0,T]$ with $t^{*}+t\leq T$ and $v_{\ker}\in\mathcal{V}_{\ker}$, this map solves the PDAE (2.1) on the time interval $[t^{*},t^{*}+t]$ with initial data $v_{\ker}+\mathcal{B}^{-}g(t^{*})$. The outcome is then the solution at time $t^{*}+t$ restricted to the kernel of $\mathcal{B}$, resulting in $u_{\ker}(t^{n+1})$, i.e., the solution at time $t^{n+1}$ within the kernel of $\mathcal{B}$,
@@ 2.3. Deterministic and probabilistic setup
By $\Psi_{\tau}$ we denote a deterministic time integration method which can be used to approximate the solution of the semi-explicit system (2.1) or its semi-discretized counterpart. Based on such a scheme, we define a randomized integrator, resulting in a sequence $\{U^{n}\}_{n\in[N]_{0}}$. Each approximation can again be decomposed into
$U^{n}=U^{n}_{\ker}+U^{n}_{c},\qquad U^{n}_{\ker}\in\mathcal{V}_{\ker},\ U^{n}_{c}\in\mathcal{V}_{\text{c}},$
with $U^{n}_{c}=\mathcal{B}^{-}g(t^{n})$. As for the continuous solution, we assume that the approximation at time $t^{n+1}$ is fully determined by $U^{n}_{\ker}$ and the right-hand sides. Hence, we write
@@
where each $\xi^{n}$ is assumed to be an i.i.d. centered Gaussian random variable with values in $\mathcal{V}_{\ker}$. Since the perturbation only acts on the dynamic part of the solution, the constraint $\mathcal{B}U^{n}=g(t^{n})$ is maintained for all $n$. Hence, we have $u_{c}(t^{n})=U^{n}_{c}$ and the sequence of errors $e^{n}$ of the random approximation is defined by
@@
Similarly to [CGS+17], the assumptions on the deterministic scheme $\Psi_{\tau}$ are summarized in the following.
@@
Let $\chi^{n}(s)$ be a centered Gaussian stochastic process taking values in $\mathcal{V}_{\ker}$, where $\chi^{n}(s)\sim\mathcal{N}(0,\sigma^{2})$ for all $s\in[0,t]$ with $\sigma\in\mathbb{R}$. The perturbation is given by $\xi^{n}(t)=\int_{0}^{t}\chi^{n}(s)\,\mathrm{d}s$ with $\chi^{n}$ being i.i.d. over different time steps and admits parameters $C_{\xi},p>0$ such that for all $t\in[0,\tau]$ it holds that $\mathbb{E}\,\|\xi^{n}(t)\|_{\mathcal{H}}^{2}\leq C_{\xi}\,t^{2p+1}$.
In line with the work [CGS+17], within the upcoming convergence analysis, we will observe that choosing $p=q$ preserves the strong order of accuracy of the underlying deterministic integrator. This choice introduces the maximum amount of noise consistent with the accuracy of the original deterministic integrator.
@@ 3. Mean Square Convergence Analysis
for all $n\in[N]$. Therein, $r$ is the mean square order of convergence of the method. For the unconstrained case, cf. [CGS+17], it is known that $r$ is equal to the minimum of $q$ (convergence order of the deterministic time integrator) and $p$ (order of the random noise). Given that our approach is based on a construction in the kernel of the constraint, one can hope to achieve the same order for the given setting. The analysis of this section will make use of the following inequality.
Lemma 3.1 (Discrete Gronwall lemma [LSS22, App. C]).
@@
In the spirit of [CGS+17, Th. 2.2], we can now state the global mean square convergence result for probabilistic time integrators applied to (2.1).
@@
where the second part of Assumption 2.7 was applied in the last step. Now we consider the expectation of this estimate. Since $\xi^{n}$ equals a zero-mean Gaussian random variable, the expectation of the inner product vanishes and we get together with Assumption 2.8
@@
Finally, an application of the triangle inequality,
$\mathbb{E}\,\|e^{n}\|^{2}_{\mathcal{H}}$ $\leq 2\,\mathbb{E}\,\|\tilde{e}^{n}\|^{2}_{\mathcal{H}}+2\,\mathbb{E}\,\|e^{n}-\tilde{e}^{n}\|^{2}_{\mathcal{H}},$
and the assumption on the global error of the deterministic solver given in (2.7) yields the desired result. ∎
@@
Analogously to [CGS+17], the convergence result of Theorem 3.2 implies that a reasonable option for the noise scale $p$ is to set $p=q$, where $q$ is the order of the underlying deterministic method. This choice of $p$ preserves the convergence of the underlying deterministic method, while providing a probabilistic interpretation of the numerical solution.
@@ 4.1. Probabilistic implicit Euler scheme
(4.1b) $\mathcal{B}u^{n+1}\hskip 80.52135pt$ $=\phantom{\tau}g(t^{n+1})\hskip 28.45274pt\qquad\text{in }\mathcal{Q}^{*}.$
Before incorporating perturbations, we discuss Assumption 2.7. It is well-known that the implicit Euler scheme converges with order one without constraints [LO93]. This then directly translates to the solution part in the kernel of the constraint, i.e., we get (2.7) with $q=1$. To prove the Lipschitz property (2.8), we use once more the decomposition of the iterates $u^{n}=u^{n}_{\ker}+u^{n}_{c}$ with $u^{n}_{c}=\mathcal{B}^{-}g(t^{n})$ as described in Section 2.2. With this, equation (4.1a) reads
$u^{n+1}_{\ker}+\mathcal{B}^{-}g(t^{n+1})$ $=u^{n}_{\ker}+\mathcal{B}^{-}g(t^{n})-\tau\,\mathcal{A}\,\big{(}u^{n+1}_{\ker}+\mathcal{B}^{-}g(t^{n+1})\big{)}$
$\hskip 59.75095pt+\tau\mathcal{B}^{*}\lambda^{n+1}+\tau\,f(t^{n+1},u_{\ker}^{n+1}+\mathcal{B}^{-}g(t^{n+1}))\quad\text{in }\mathcal{V}^{*}.$
Restricting this equation to test functions in $\mathcal{V}_{\ker}$, the Lagrange multiplier disappears. Moreover, setting $\mathcal{V}_{\text{c}}$ to the annihilator of $\mathcal{V}_{\ker}$, also $\mathcal{A}\mathcal{B}^{-}g(t^{n+1})$ vanishes and we get
$u^{n+1}_{\ker}=u^{n}_{\ker}+\tau\,\big{(}-\mathcal{A}\,u^{n+1}_{\ker}+f(t^{n+1},u_{\ker}^{n+1}+\mathcal{B}^{-}g(t^{n+1}))\big{)}-\mathcal{B}^{-}\big{(}g(t^{n+1})-g(t^{n})\big{)}$
in $\mathcal{V}_{\ker}^{*}$ and, hence,
@@
(4.2) $\begin{bmatrix}\operatorname{id}+\tau\mathcal{A}&\tau\mathcal{B}^{*}\\ \mathcal{B}&0\end{bmatrix}\begin{bmatrix}U^{n+1}\\ \lambda^{n+1}\end{bmatrix}=\begin{bmatrix}U^{n}+\tau f(t^{n+1},U^{n+1})+\xi^{n}(\tau)\\ g(t^{n+1})\end{bmatrix},$
where $\xi^{n}(\tau)\in\mathcal{V}_{\ker}$ can be defined as $\sigma\tau^{3/2}\tilde{\xi}^{n}$ by the natural choice of $p=q=1$. Here, $\sigma$ is a constant noise scale and $\tilde{\xi}^{n}$ is defined as a random vector where each component is an i.i.d. standard normal random variable, i.e., $\tilde{\xi}_{i}^{n}\sim\mathcal{N}(0,1)$, with $i$ indexing the components of the vector. The existence of a unique solution to (4.2) is guaranteed by [AZ20, Lem. 3.1]. In Section 5, we discuss the case where $\sigma$ is chosen problem-dependent such that the expected order of convergence is maintained.
@@
Note that the perturbation is applied in a way that maintains the structure of the problem, ensuring that the conditions for solvability, such as the ellipticity of $\mathcal{A}$ on the kernel of $\mathcal{B}$ and the inf–sup condition for $\mathcal{B}$, remain valid. This guarantees that adding perturbations does not affect the existence or uniqueness of the solution.
Remark 4.4.